OpenStudy (bohotness):

help please Samantha's printer can print 21 pages in 4 minutes. At that rate, how long would it take for Samantha's printer to print 105 pages? 420 minutes 16 minutes 20 minutes 26 minutes willl medal and fan

OpenStudy (anonymous):

20 minutes, just multiply 21 by x until you find a number that works for 105. however many times you multiply 21 by, you multiply the same by the time.

OpenStudy (bohotness):

oo i throughgt i had to add them all up

OpenStudy (bohotness):

Jeff, Jose, Mandy, and Whitney are participants in the school home run contest. Participant Home runs hit Home runs attempted Jeff 20 out of 30 Jose 8 out of 12 Mandy 12 out of 16 Whitney 14 out of 21 Using the scores shown, which participant has a ratio of home runs hit to home runs attempted that is not proportional to any other? Mandy Whitney Jose Jeff i think its c or b

OpenStudy (anonymous):

it is Jose

OpenStudy (bohotness):

it was worng

OpenStudy (jhannybean):

for your first question, I would solve it like this,:\[\sf\frac{21 ~\text{pages}}{4~ \text{min}} =\frac{105~ \text{pages}}{x~\text{min}}\]

OpenStudy (bohotness):

?

OpenStudy (jhannybean):

The question you posted to open this thread? :P Im just showing visually how I would go about solving it in ratio proportions :)

OpenStudy (bohotness):

oo ok

OpenStudy (bohotness):

i dont understand this one radio station gave away 228 tickets to a concert. They gave away 5 times as many tickets to listeners as they gave to employees. How many tickets did they give away to employees? 117 38 222 190 i got 875 because i times 5 times 5 times 5 time 5 times 5

OpenStudy (anonymous):

divide by five, then find the difference

OpenStudy (anonymous):

medal?? @bohotness

OpenStudy (bohotness):

thats 25 then

OpenStudy (bohotness):

ok

OpenStudy (jhannybean):

Oh oops. I miscounted # of tickets.

OpenStudy (bohotness):

ok

OpenStudy (jhannybean):

228 = total # of tickets x = # of tickets given to employees 5x = # of tickets given to listeners 5x + x = 228 6x = 228 x = 228/6

OpenStudy (bohotness):

ooo okay

OpenStudy (bohotness):

While Joey was standing outside on a sunny day, his shadow was 10 feet long. The shadow of a nearby Oak tree was 15 feet long. If Joey is 6 feet tall, how tall is the Oak tree? 9.6 feet 6.5 feet 25 feet 9 feet 0 and thi got 13e another one i got the answer now

OpenStudy (anonymous):

10 times 15 divided by 6 @bohotness